. Find the square of the following numbers. i. 32 ii. 35 iii. 86 iv. 93 v. 71 vi. 46 
 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

i. (32)2
= (30 +2)2
= (30)2+ (2)2+ 2×30×2 =900+4+120

= 1024

ii. (35)2
= (30+5 )2
= (30)2+ (5)2+ 2×30×5 =900+25+300
= 1225

iii. (86)2
=(90-4)2
= (90)2+ (4)2- 2×90×4 =8100+16–720 =8116–720
= 7396

iv. (93)2
= (90+3 )2
= (90)2+ (3)2+ 2×90×3 =8100+9+540
= 8649

v. (71)2
= (70+1 )2
= (70)2+ (1)2+2×70×1 =4900+1+140
= 5041

vi. (46)2
=(50-4)2
= (50)2+ (4)2- 2×50×4 =2500+16–400
= 2116


Thanks